The sum of two numbers is 27.Half the smaller number plus twice the larger number is 36.What are the numbers
Answer by mananth(16949)

let the number be x,
the other number will be 27-x.
.
half the smaller number = x/2
twice the larger number = 2(27-x)
.
x/2 + 2(27-x)= 36
x/2 +54 -2x = 36
LCM =2
x+108 -4x= 2*36
-3x=-108+72
-3x=-36
x= 12
27-12 =15
12, 15 are the numbers